# Verifying Well-Posedness of Linear PDEs using Convex Optimization

### 01 Apr 2026 v1

@@ VI Examples
Given the well-posedness test for linear PDEs in Cor. 13, we now verify the efficacy of this test for verifying well-posedness of several classical PDE examples, as well as a more complicated system. For each example, the optimization program in (6) is implemented in PIETOOLS [16], with $\epsilon_{1}=\epsilon_{2}=10^{-1}$, and the resulting semidefinite program is solved using Mosek [20]. For each of the classical PDE examples, an ill-posed version is also formulated by modifying either the definition of $A$ or of its domain, and we verify that the LPI in (6) becomes infeasible. Analytic proofs of well-posedness for the examples are provided in Appx. A.
@@ VI-A1 Well-Posed Example
We formulate an associated Cauchy problem as in (3) on $H:=L_{2}[0,1]$, by defining $A:D\to L_{2}[0,1]$ as
$A:=-\partial_{s},\qquad D:=\bigl\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(0)=0\bigr\}.$
Then $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $H$, with exponential growth rate $\omega$ and gain $M=e^{-\omega}$ for any $\omega\leq 0$. To verify well-posedness using Cor. 13, define PI operators $(\mathcal{T}\mathbf{v})(s):=\int_{0}^{s}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$. Then $\{\mathcal{T},\mathcal{A}\}$ define the equivalent PIE representation, with $\mathcal{T}^{-1}=\partial_{s}$ and $\mathcal{A}=A\circ\mathcal{T}$. Solving the optimization program in (6) with $\mathcal{P}=I$, well-posedness of the PDE can be verified with $\omega=0$, certifying that $A$ generates a contraction semigroup. Furthermore, parameterizing a more general PI operator variable $\mathcal{P}$ using monomials of degree at most $2d$, a least upper bound on the exponential growth rate $\omega$ of the semigroups is established as in Table I.
@@
VI-A2 Ill-Posed Example
Consider now the transport equation with boundary condition applied at the opposite boundary,
$u_{t}(t,s)=-u_{s}(t,s),\qquad u(t,1)=0.$
In this case the PDE is not well-posed, admitting no solution for nontrivial initial states. As such, defining the associated PIE representing, setting $(\mathcal{T}\mathbf{v})(s)=-\int_{s}^{1}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
@@ VI-B1 Well-Posed Example
Consider the reaction–diffusion equation
@@
and $\mathcal{A}:=A\circ\mathcal{T}=I+r\mathcal{T}$. Solving the LPI in (6), parameterizing $\mathcal{P}$ by polynomials of degree at most $6$, well-posedness can be verified for any $r\in\{0,1,\dots,20\}$, with least upper bound $\omega=\pi^{2}-r-10^{-5}$ on the exponential growth rate.
@@ VI-B2 Ill-Posed Example
Consider the reverse-time heat equation
$u_{t}(t,s)$ $=-u_{ss}(t,s),$ $t\geq 0,\ s\in[0,1],$
$u(t,0)$ $=u(t,1)=0.$
Defining $A=-\partial_{s}^{2}$ on the same domain $D$ as before, the resulting Cauchy problem is not well-posed. Consequently, defining the associated PIE representation by $\mathcal{T}$ as in (8) and $\mathcal{A}=-I$, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
@@ VI-C1 Well-Posed Example
$u_{t}(t,0)$ $=u_{s}(t,1)=0.$ (10)
Defining $\mathbf{u}(t)=(u_{t}(t),u_{s}(t))$, this PDE can be written as in (3) with
@@ VI-C2 Ill-Posed Example
For the same wave equation in (9), suppose we instead define the state $\mathbf{u}(t)=(u(t),u_{t}(t))$ with
@@
Then, the resulting Cauchy problem is not well-posed. Computing $\mathcal{T}=(\partial_{s}^{2})^{-1}$ and $\mathcal{A}=A\circ\mathcal{T}$ yields a PIE representation whose operator $\mathcal{Q}:=(\lambda\mathcal{T}-\mathcal{A})(\lambda\mathcal{T}-\mathcal{A})^{*}$ cannot satisfy $\mathcal{Q}\succeq\epsilon I$ for any $\epsilon>0$ or $\lambda\in\mathbb{R}$. Consequently, the LPI (6) is infeasible for all $\omega\in\mathbb{R}$.
@@ VI-D Coupled PDEs
To illustrate the application of Cor. 13 for verifying well-posedness of less classically well-studied PDEs, consider finally the following system of a coupled PDEs
@@
Introducing the state $\mathbf{u}(t)=(u_{1}(t),u_{2}(t),\partial_{s}u_{3}(t),\partial_{t}u_{3}(t))$, this system can be formulated as in (3) using a suitable operator $A:D\to L_{2}^{4}[0,1]$ on the domain
$D:=\biggl\{\mathbf{u}\in{\small\left[\!\begin{array}[]{l}W_{2}^{1}[0,1]\\ W_{2}^{2}[0,1]\\ W_{2}^{1,2}[0,1]\end{array}\!\right]}\,\Bigl|\ {\small\begin{array}[]{l}\mathbf{u}_{1}(0)=\mathbf{u}_{2}(1)=\mathbf{u}_{3}(0)\\ \mathbf{u}_{2}(0)=\mathbf{u}_{4}(1)=0\\ \partial_{s}\mathbf{u}_{2}(1)=\mathbf{u}_{4}(0)\end{array}}\biggr\}$
@@ VII Conclusion
A PIE-based framework for verifying well-posedness of partial differential equations was developed. By reformulating the Lumer–Phillips generation theorem in terms of Partial Integral (PI) operators, necessary and sufficient conditions were established for a PIE to generate a quasicontraction semigroup. These conditions were relaxed to determine sufficient conditions for generation of a more general $C_{0}$-semigroup, with upper bound on the exponential growth rate of solutions. Using these results, a convex optimization program on PI operator variables was developed for testing well-posedness of any PDE admitting an equivalent PIE representation. The resulting conditions can be implemented using PIETOOLS and solved using semidefinite programming. Numerical examples demonstrated that the method can verify well-posedness and accurately estimate growth rates for several 1D PDE systems.
@@ A-A Transport Equation
Consider first the operator
$A:=-\partial_{s}:D\to L_{2}[0,1],$
@@
defining a transport equation with Dirichlet boundary condition at the lower boundary. A solution to the transport equation can be readily found using the method of characteristics, noting that $u_{t}(t,s)=-u_{s}(t,s)$ and $u(0,s)=\mathbf{x}(s)$ implies $u(t,s)=\mathbf{x}(s-t)$ for all $t\leq s$. Imposing the boundary condition, $u(t,0)$, it follows that $A$ is the infinitesimal generator of the following $C_{0}$-semigroup.
@@
Here, $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq\left\lVert{e^{t^{\prime}A}\mathbf{x}}\right\rVert_{L_{2}}$ for all $t\geq t^{\prime}\geq 0$ and $\mathbf{x}\in D$, whence $(e^{tA})_{t\geq 0}$ is a contraction semigroup. Furthermore since $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}=0$ for all $t\geq 1$ and $\mathbf{x}\in D$, it follows that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq e^{\omega[t-1]}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for any $\omega\in\mathbb{R}$ all $t\geq 1$ and $\mathbf{x}\in D$. Thus, the $C_{0}$-semigroup generated by $A$ has exponential growth rate $\omega$ for any $\omega\in\mathbb{R}$, with gain $M=e^{\omega}$.
However, consider now the following operator,
$A:=-\partial_{s}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(1)=0\},$
imposing a Dirichlet boundary condition on the upper boundary. In this case, the semigroup generated by $A$ would still need to satisfy $(e^{tA}\mathbf{x})(s)=\mathbf{x}(t-s)$ for $t\leq s$. However, for any nontrivial $\mathbf{x}\in D$, we may not have $\mathbf{x}(t-1)=0$ for arbitrary $t\geq 0$, implying that $e^{tA}\mathbf{x}\notin D$ for all $t\geq 0$. Therefore, $A$ cannot be the infinitesimal generator of a $C_{0}$-semigroup, implying the associated transport equation is ill-posed.
@@ A-B Heat Equation
Next, consider the operator
$A:=\partial_{s}^{2}:D\to L_{2}[0,1],$
@@
defining the heat equation with Dirichlet boundary conditions. The solution to the heat equation can be readily established using separation of variables, finding that $A$ generates the $C_{0}$-semigroup
$(e^{At}\mathbf{x})(s)=\sum_{n=1}^{\infty}2\left\langle\mathbf{x},\boldsymbol{\phi}_{n}\right\rangle_{L_{2}}e^{-n^{2}\pi^{2}t}\boldsymbol{\phi}_{n}(s),$
where $\boldsymbol{\phi}_{n}(s):=\sin(n\pi s)$ for $n\in\mathbb{N}$. However, consider now the reverse-time heat equation, defined by
$A:=-\partial_{s}^{2}:D\to L_{2}[0,1],$
$D:=\{\mathbf{u}\in W_{2}^{2}[0,1]\mid\mathbf{u}(0)=\mathbf{u}(1)=0\}.$
It is clear that if this operator generates a $C_{0}$-semigroup, then this semigroup would have to be given by
$(e^{At}\mathbf{x})(s)=\sum_{n=1}^{\infty}2\left\langle\mathbf{x},\boldsymbol{\phi}_{n}\right\rangle_{L_{2}}e^{n^{2}\pi^{2}t}\boldsymbol{\phi}_{n}(s).$
However, this semigroup is not continuous, since for any $M>0$ and $t>0$, we can establish $n\in\mathbb{N}$ sufficiently large such that $e^{n\pi^{2}t}>2M$, whence $\left\lVert{e^{At}\boldsymbol{\phi}_{n}}\right\rVert_{L_{2}}>M$. Thus, the operator $A$ does not generate a $C_{0}$-semigroup, implying the associated reverse-time heat equation is not well-posed.

### 02 Apr 2026 v2

@@ VI Illustrations of the Well-Posedness Test
Given the well-posedness test for linear PDEs in Cor. 13, we now verify the effectiveness of this test for testing well-posedness of several classical PDE examples. We then illustrate the utility of the test by applying it to a more complicated, speculative PDE for which well-posedness has not been previously studied. For each example, the optimization program in (7) is implemented in PIETOOLS [16], with $\epsilon_{1}=\epsilon_{2}=10^{-1}$ and $\lambda=\omega+1$, and the resulting semidefinite program is solved using Mosek [21]. For each of the classical PDE examples, an ill-posed version is also formulated by modifying either the definition of $A$ or of its domain, and we verify that the LPI in (7) becomes infeasible. Although well-posedness results for the classical PDE examples are well-established, an analytic proof of well-posedness is also provided in Appx. A.
@@ VI-A1 Well-Posed Example
We formulate this PDE as in (3), by defining $A:=-\partial_{s}:D\to L_{2}[0,1]$ for $D:=\bigl\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(0)=0\bigr\}$. Then $A$ is the infinitesimal generator of a $C_{0}$-semigroup on $L_{2}$, with exponential growth rate $\omega$ and gain $M=e^{-\omega}$ for any $\omega\leq 0$. We verify this using Cor. 13, defining the associated PIE representation by $(\mathcal{T}\mathbf{v})(s):=\int_{0}^{s}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$. Solving the optimization program in (7) with $\mathcal{P}=I$, well-posedness can be verified with $\omega=0$, certifying that $A$ generates a contraction semigroup. Furthermore, parameterizing a more general PI operator variable $\mathcal{P}$ using monomials of degree at most $2d$, an improved least upper bound on the exponential growth rate $\omega$ of solutions is established as in Table I.
VI-A2 Ill-Posed Example
Consider now the transport equation with Dirichlet condition at the opposite boundary, so that $A:=-\partial_{s}$ and $D:=\bigl\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(1)=0\bigr\}$, which is ill-posed. Defining the associated PIE representation – setting $(\mathcal{T}\mathbf{v})(s)=-\int_{s}^{1}\mathbf{v}(\theta)d\theta$ and $\mathcal{A}:=-I$ – the LPI (7) is infeasible for all $\omega\in\mathbb{R}$. We can verify this using $\mathbf{v}_{n}\in L_{2}[0,1]$ defined by $\mathbf{v}_{n}(s):=n$ for $s\in[0,\frac{1}{n}]$ and $\mathbf{v}_{n}(s):=0$ else, for $n\in\mathbb{N}$. Then
$(\mathcal{T}\mathbf{v}_{n})(s)=\begin{cases}ns-1,&s<\frac{1}{n},\\ 0,&\textnormal{else},\end{cases}\qquad n\in\mathbb{N},$
so that $\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{T}\mathbf{v}_{n}\right\rangle_{L_{2}}=\frac{1}{3n}$ and $\left\langle\mathcal{T}\mathbf{v}_{n},\mathbf{v}_{n}\right\rangle=-\frac{1}{2}$. It follows that, for any $\omega\in\mathbb{R}$ and $\mathcal{P}\in\mathcal{L}(L_{2})$ satisfying $\mathcal{P}\succeq\epsilon_{1}^{2}I$, letting $n\to\infty$,
$\frac{\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{A}\mathbf{v}_{n}\right\rangle_{L_{2,\mathcal{P}}}}{\omega\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{T}\mathbf{v}_{n}\right\rangle_{L_{2,\mathcal{P}}}}\geq\frac{\epsilon_{1}^{2}}{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}\omega}\frac{-\left\langle\mathcal{T}\mathbf{v}_{n},\mathbf{v}_{n}\right\rangle_{L_{2}}}{\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{T}\mathbf{v}_{n}\right\rangle_{L_{2}}}\to\infty,$
whence Condition 1) of Thm. 12 cannot be satisfied.
@@ VI-B1 Well-Posed Example
Next, consider the following reaction–diffusion equation,
@@
and $\mathcal{A}:=A\circ\mathcal{T}=I+r\mathcal{T}$. Solving the LPI in (7), parameterizing $\mathcal{P}$ by polynomials of degree at most $6$, well-posedness can be verified for any $r\in\{0,1,\dots,20\}$, with least upper bound $\omega=r-\pi^{2}+10^{-5}$ on the exponential growth rate.
@@ VI-B2 Ill-Posed Example
Consider now the following ill-posed, reverse-time heat equation,
$A:=-\partial_{s}^{2},\quad D=\bigl\{\mathbf{u}\in W_{2}^{2}[0,1]\mid\mathbf{u}(0)=\mathbf{u}(1)=0\bigr\},$
The associated PIE representation is defined by $\mathcal{T}$ as in (9) and $\mathcal{A}=-I$, where we remark that $\mathcal{T}=$-$\mathcal{R}^{*}\mathcal{R}$ for $\mathcal{R}:=\partial_{s}\circ\mathcal{T}$. Considering then $\mathbf{v}_{n}(s):=-n\sin(n\pi s)$, we have for each $n\in\mathbb{N}$ that $(\mathcal{R}\mathbf{v}_{n})(s)=\cos(n\pi s)$ and $(\mathcal{T}\mathbf{v}_{n})(s)=\frac{1}{n}\sin(n\pi s)$, whence $\left\lVert{\mathcal{R}\mathbf{v}_{n}}\right\rVert_{L_{2}}^{2}=\frac{1}{2}$ and $\left\lVert{\mathcal{T}\mathbf{v}_{n}}\right\rVert_{L_{2}}^{2}=\frac{1}{2n^{2}}$. It follows that, for any $\omega\in\mathbb{R}$ and $\mathcal{P}\in\mathcal{L}(L_{2})$ with $\mathcal{P}\succeq\epsilon^{2}I$, letting $n\to\infty$,
$\frac{\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{A}\mathbf{v}_{n}\right\rangle_{L_{2,\mathcal{P}}}}{\omega\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{T}\mathbf{v}_{n}\right\rangle_{L_{2,\mathcal{P}}}}\geq\frac{\epsilon_{1}^{2}}{\left\lVert{\mathcal{P}}\right\rVert_{\mathcal{L}(L_{2})}\omega}\frac{\left\langle\mathcal{R}\mathbf{v}_{n},\mathcal{R}\mathbf{v}_{n}\right\rangle_{L_{2}}}{\left\langle\mathcal{T}\mathbf{v}_{n},\mathcal{T}\mathbf{v}_{n}\right\rangle_{L_{2}}}\to\infty.$
Thus, Condition 1) of Thm. 12 cannot be satisfied, and the well-posedness LPI (7) is infeasible.
@@ VI-C1 Well-Posed Example
$u_{t}(t,0)$ $=u_{s}(t,1)=0.$
To allow for well-posedness, we consider a representation in terms of $\mathbf{u}(t)=(\mathbf{u}_{1}(t),\mathbf{u}_{2}(t))=(u_{t}(t),u_{s}(t))$ (which does not include $u$), expressing the PDE dynamics of $\mathbf{u}$ as in (3) with
@@ VI-C2 Ill-Posed Example
For the same wave equation in (10), suppose we instead define the state $\mathbf{u}(t)=(\mathbf{u}_{1}(t),\mathbf{u}_{2}(t))=(u(t),u_{t}(t))$ with
@@
Then, the resulting PDE in (3) is not well-posed, and LPI (7) is infeasible for the associated PIE representation. In particular, computing $\mathcal{T}=(\partial_{s}^{2})^{-1}$ and $\mathcal{A}=A\circ\mathcal{T}$, the resulting operator $\mathcal{Q}:=(\lambda\mathcal{T}-\mathcal{A})(\lambda\mathcal{T}-\mathcal{A})^{*}$ for any $\lambda\in\mathbb{R}$ will be of the form
$(\mathcal{Q}\mathbf{v})(s)=\begin{bmatrix}0&0\\ 0&1\end{bmatrix}\mathbf{v}+\int_{0}^{1}\boldsymbol{K}(s,\theta)\mathbf{v}(\theta)d\theta,$
for some (semi-separable) kernel $\boldsymbol{K}$. Since the multiplier term herein acts only on $\mathbf{v}_{2}$, this operator cannot satisfy $\mathcal{Q}\succeq\epsilon I$ for any $\epsilon>0$, whence Condition 2) in Thm. 12 is infeasible.
@@ VI-D Coupled PDEs
To illustrate the application of Cor. 13 for verifying well-posedness of less classically well-studied PDEs, consider finally the following, entirely speculative, system of coupled PDEs
@@
Introducing the state $\mathbf{u}(t)=(u_{1}(t),u_{2}(t),\partial_{t}u_{3}(t),\partial_{s}u_{3}(t))$, this system can be formulated as in (3) using a suitable operator $A:D\to L_{2}^{4}[0,1]$ on the domain
$D:=\biggl\{\mathbf{u}\in{\small\left[\!\begin{array}[]{l}W_{2}^{1}[0,1]\\ W_{2}^{2}[0,1]\\ W_{2}^{1,2}[0,1]\end{array}\!\right]}\,\Bigl|\ {\small\begin{array}[]{l}\mathbf{u}_{1}(0)=\mathbf{u}_{2}(1)=\mathbf{u}_{3}(0)\\ \mathbf{u}_{2}(0)=\mathbf{u}_{4}(1)=0\\ \partial_{s}\mathbf{u}_{2}(1)=\mathbf{u}_{4}(0)\end{array}}\biggr\}.$
@@ VII Conclusion
A PIE-based framework for verifying well-posedness of partial differential equations was developed. By reformulating the Lumer–Phillips generation theorem in terms of Partial Integral (PI) operators, necessary and sufficient conditions were established for a PIE to generate a quasicontraction semigroup. These conditions were relaxed to determine sufficient conditions for generation of a more general $C_{0}$-semigroup, with upper bound on the exponential growth rate of solutions. Using these results, a convex optimization program on PI operator variables was developed for testing well-posedness of any PDE admitting an equivalent PIE representation. The resulting conditions can be implemented using PIETOOLS and solved using semidefinite programming. Illustrative examples demonstrated that the method can verify well-posedness and accurately estimate growth rates for several 1D PDE systems.
@@ A-A Transport Equation
Consider first $A:=-\partial_{s}:D\to L_{2}[0,1]$ on
@@
defining a transport equation with Dirichlet boundary condition at the lower boundary. A solution to the transport equation can be readily found using the method of characteristics, noting that $u_{t}(t,s)=-u_{s}(t,s)$ and $u(0,s)=\mathbf{x}(s)$ imply $u(t,s)=\mathbf{x}(s-t)$ for all $t\leq s$. Imposing $u(t,0)=0$, it follows that $A$ generates the $C_{0}$-semigroup
@@
Here, $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq\left\lVert{e^{\tau A}\mathbf{x}}\right\rVert_{L_{2}}$ for all $t\geq\tau\geq 0$ and $\mathbf{x}\in D$, whence $(e^{tA})_{t\geq 0}$ is a contraction semigroup. Furthermore since $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}=0$ for all $t\geq 1$ and $\mathbf{x}\in D$, it follows that $\left\lVert{e^{tA}\mathbf{x}}\right\rVert_{L_{2}}\leq e^{\omega[t-1]}\left\lVert{\mathbf{x}}\right\rVert_{L_{2}}$ for any $\omega\in\mathbb{R}$ and all $t\geq 1$ and $\mathbf{x}\in D$. Thus, the $C_{0}$-semigroup generated by $A$ has exponential growth rate $\omega$ for any $\omega\in\mathbb{R}$, with gain $M=e^{\omega}$.
However, consider now $A:=-\partial_{s}:\hat{D}\to L_{2}[0,1]$ on $\hat{D}:=\{\mathbf{u}\in W_{2}^{1}[0,1]\mid\mathbf{u}(1)=0\}$, imposing a Dirichlet boundary condition on the upper boundary. In this case, a $C_{0}$-semigroup generated by $A$ would still need to satisfy $(e^{tA}\mathbf{x})(s)=\mathbf{x}(t-s)$ for $t\leq s$. However, for nontrivial $\mathbf{x}\in D$, $\mathbf{x}(t-1)\neq 0$ for arbitrary $t\geq 0$, implying that $e^{tA}\mathbf{x}\notin D$ for all $t\geq 0$. Therefore, the associated transport equation is ill-posed.
@@ A-B Heat Equation
Next, consider $A:=\partial_{s}^{2}:D\to L_{2}[0,1]$ on
@@
defining the heat equation with Dirichlet boundary conditions. The unique solution to the heat equation can be readily established using separation of variables, finding that $A$ generates the $C_{0}$-semigroup with growth rate $\omega=-\pi^{2}$,
$(e^{tA}\mathbf{x})(s)=\sum_{n=1}^{\infty}2\left\langle\mathbf{x},\boldsymbol{\phi}_{n}\right\rangle_{L_{2}}e^{-n^{2}\pi^{2}t}\boldsymbol{\phi}_{n}(s),$
where $\boldsymbol{\phi}_{n}(s):=\sin(n\pi s)$ for $n\in\mathbb{N}$. Consider now the reverse-time heat equation, defined by $\hat{A}:=-\partial_{s}^{2}:D\to L_{2}[0,1]$, on the same domain $D$. Then the $C_{0}$-semigroup generated by $\hat{A}$ would have to satisfy $e^{t\hat{A}}=e^{-tA}$. However, this semigroup is not continuous, since for any $M>0$ and $t>0$, we can establish $n\in\mathbb{N}$ sufficiently large such that $e^{n\pi^{2}t}>2M$, whence $\left\lVert{e^{t\hat{A}}\boldsymbol{\phi}_{n}}\right\rVert_{L_{2}}>M$. Thus, $\hat{A}:D\to L_{2}[0,1]$ does not generate a $C_{0}$-semigroup, implying the reverse-time heat equation is not well-posed.